# Study of transport properties of a hot and dense QCD matter using a novel approximation method

### v1

@@ 4 Applications
The purpose of this section is to study some applications of electrical and thermal conductivities for the hot and dense QCD medium. In subsection 4.1, we analyze the heat transfer in the medium through the thermal diffusion constant in the novel RTA and the novel BGK models. The relative behavior of the thermal conductivity and the electrical conductivity is explored in subsection 4.2 using the Lorenz number in the aforesaid models.
@@ 4.1 Thermal diffusion constant
Figure 5 shows the variation of the thermal diffusion constant with the temperature at a fixed chemical potential in the novel RTA and the novel BGK models. It is evident from figures 5(a) and 5(b) that the magnitude of gets decreased regardless of the relaxation type ( or ) for both RTA and BGK models with the modified integral term. When we compare it with the standard RTA and BGK models, we have found that the momentum-dependent relaxation time () leads to a considerable shift in the thermal diffusion constant. We can conclude from the thermal diffusion constant related ratios in tables 1 and 2 that, for the novel RTA and the novel BGK models, the decline in the magnitude of becomes larger when considering the momentum-dependent relaxation time as compared to the standard RTA and BGK models.
@@ 5 Summary
In this work, we studied the transport properties of a hot and dense QCD matter in the novel RTA and the novel BGK models. This study could help to get an understanding of the charge and heat transport phenomena in a strongly coupled thermal QCD medium with a finite chemical potential background through the novel approaches of the aforesaid models. Using the kinetic theory technique, we solved the relativistic Boltzmann transport equation with a modified collision integral for estimating the thermal and electrical conductivities of the aforementioned medium. We examined the effects of the modified collision integral on the charge and heat transport in a thermal QCD medium, in contrast to the widely used collision terms of the RTA and the BGK models. In this research, we used the quasiparticle description of partons, where both the temperature and the chemical potential affect the constituent masses in the medium. Our modifications to the collision integral for the RTA and the BGK models were shown to result in a considerable drop in the electrical and thermal conductivities. Additionally, we found that these transport coefficients exhibit a discernible shift when we take into account a momentum-dependent relaxation time (). In the same regime, the thermal diffusion constant and the Lorenz number were also investigated. We observed that the novel collision integral leads to an increase in the Lorenz number and a decrease in the thermal diffusion constant of the medium.
@@ 6 Acknowledgments
One of us (S. R.) is thankful to the Indian Institute of Technology Bombay for the Institute postdoctoral fellowship and S. D. acknowledges the SERB Power Fellowship, SPF/2022/000014 for the support on this work.
@@ Appendix A Derivation of equation (20)
Similarly, the 3rd and 4th terms can be respectively reduced to
@@
Appendix B Derivation of equation (28)
@@
For a proper treatment of the right-hand side of the above equation, we are considering an infinitesimal perturbation to the equilibrium distribution function, i.e. with and after linearizing it, we get
Now, incorporating the result of ref. [65] in the novel BGK model, we get
Using eq. (B), eq. (B.69) and the partial derivatives as given in eq. (A.66), eq. (B.67) takes the following form,
After some mathematical calculation, we get
where is related to the infinitesimal change of the distribution function () by the relation given in eq. (26). After putting the value of in eq. (26), we get as given in eq. (28).
Appendix C Derivation of equation (46)
@@ Appendix D Derivation of equation (52)
With the help of eq. (B), eq. (B.69) and the partial derivatives as given in eq. (C.73), we can express the above equation in the novel BGK model as

### v3

@@ 4 Applications
The purpose of this section is to study some applications of the electrical and the thermal conductivities for the hot and dense QCD medium. In subsection 4.1, we analyze the rate of heat transfer in the medium through the thermal diffusion constant in the novel RTA and the novel BGK models. The relative behavior of the thermal conductivity and the electrical conductivity is explored in subsection 4.2 using the Lorenz number in the aforesaid models.
@@ 4.1 Thermal diffusion constant
Figure 5 shows the variation of the thermal diffusion constant with the temperature at a fixed chemical potential in the novel RTA and the novel BGK models. It is evident from figures 5(a) and 5(b) that the magnitude of gets decreased regardless of the relaxation type ( or ) for both RTA and BGK models with the modified integral term. When we compare it with the standard RTA and standard BGK models, we have found that the momentum-dependent relaxation time () leads to a considerable shift in the thermal diffusion constant. We can conclude from the thermal diffusion constant related ratios in tables 1 and 2 that, for the novel RTA and the novel BGK models, the decline in the magnitude of becomes larger when considering the momentum-dependent relaxation time as compared to the standard RTA and standard BGK models.
@@ 5 Summary
In this work, we studied the transport properties of a hot and dense QCD matter in the novel RTA and the novel BGK models. From this study, we gained more accurate understanding of the charge and heat transport phenomena in a strongly interacting thermal QCD medium with a finite chemical potential background by using the novel approaches of the aforesaid models. Using the kinetic theory technique, we solved the relativistic Boltzmann transport equation with a modified collision integral for estimating the thermal and electrical conductivities of the medium. We examined the effects of the modified collision integral on the transport of charge and heat in a thermal QCD medium, in contrast to the widely used collision terms of the RTA and the BGK models. In this research, we used the quasiparticle description of partons, where both the temperature and the chemical potential affect the constituent masses in the medium. By introducing a novel collision integral, we demonstrated that the charge and heat conductivities in QGP got significantly lower than the previously estimated results, which has direct implications for electromagnetic radiation from QGP and hydrodynamic modeling in heavy ion collisions, indicating reduced soft photon and dilepton production rates in RHIC and LHC experiments. Our findings contribute to a more accurate theoretical framework for understanding the evolution of QGP in RHIC and LHC experiments. Additionally, when we replaced the momentum-independent relaxation time () with the momentum-dependent relaxation time (), we observed a discernible shift in these transport coefficients, thus explaining the impact of the momentum-dependent relaxation time on the transport properties of the medium. In the same regime, the thermal diffusion constant and the Lorenz number were also investigated. We observed that the novel collision integral leads to an increase in the Lorenz number and a decrease in the thermal diffusion constant of the medium.
This study provides new insights into charge and heat transport phenomena in a strongly interacting QCD medium. Our findings have phenomenological implications for measurable quantities in heavy ion collisions. The electrical conductivity governs the interaction of QGP with electromagnetic fields, and it is related to the photon and dilepton generation from QGP via the electromagnetic current. Reduced electrical conductivity in the novel RTA and the novel BGK models as identified in our work leads to reduced photon and dilepton emission rates, hence influencing the electromagnetic signatures of the quark-gluon plasma as well as the existing experimental assessments. Reduced soft photon and dilepton yields have been seen at RHIC and LHC, suggesting that the charge transport in the quark-gluon plasma is more constrained than that predicted by standard models. Our study suggests that the QGP charge transport is more suppressed than in the traditional kinetic models, aligning with these experimental observations. Lower value of the thermal conductivity in the novel approaches of the RTA and the BGK models implies that the temperature variations persist for longer durations, leading to localized hotspots within the quark-gluon plasma. The suppression of the heat diffusion in the novel approaches of the RTA and the BGK models in our work is consistent with the strong local temperature gradients as seen at RHIC experimental facility [94, 95, 96]. Our finding regarding the increased Lorenz number in the novel RTA and the novel BGK models suggests that the QGP behaves as a strongly coupled system, where the charge transport is more suppressed than the heat transport, when compared to the predictions of the standard RTA and standard BGK models. Although our novel collision integral offers a more precise characterization of QGP transport properties, enhancements might be achieved by including the nonequilibrium corrections and higher-order moments in the Boltzmann transport equation. Subsequent research should investigate anisotropic transport coefficients in the nonequilibrium quark-gluon plasma configurations. Indirect testing of our results could be derived from experimental measures of soft photon generation, elliptic flow and jet quenching.
@@ 6 Acknowledgments
One of the present authors (S. R.) acknowledges financial support from ANID Fondecyt Postdoctoral Grant 3240349 and S. D. acknowledges the SERB Power Fellowship, SPF/2022/000014 for the support on this work.
@@ Appendix A Derivation of equation (20)
Here, we have used the fact that the shift of the distribution function from its equilibrium value is infinitesimal, thus one can use the derivative . Similarly, the third and the fourth terms can be respectively reduced to
@@
Appendix B The particle number and the energy-momentum conservation laws
In this study, we have introduced a novel collision integral in the BGK model of the following form,
We will be subject to the following constraints if we insist that the conservation laws be upheld.
Validity of the particle number conservation law in the novel BGK model:
where,
Now, putting the results of equations (B.77), (B.78) and (B.79) in eq. (B.76), we get
From eq. (B.80), it is confirmed that the particle number is instantaneously conserved in the novel BGK model.
Validity of the energy-momentum conservation law in the novel BGK model:
where , and are respectively expressed as
The energy-momentum conservation requires that,
The term can be calculated as follows.
where,
Now, the term turns out to be
In getting the above equation, we have used the following substitutions (in this case, is expressed in terms of an equilibrium contribution and a nonequilibrium correction () as ),
The term can be calculated as follows.
where,
Thus the term becomes
We have obtained the above equation by using the following substitutions,
The term can be calculated as follows.
Now, substituting the results of equations (B.91), (B.101) and (B.105) in eq. (B), we get
The above equation serves as the matching condition for the novel BGK model and can be expressed as
where,
In the present scenario, Q1 and Q2 represent the constraints for the conservation of the particle number and the energy-momentum, respectively. The consistency of our formulation within the generalised hydrodynamic frame described in ref. [72] may be readily observed from equations (B.108) and (B.109).
Appendix C Derivation of equation (34)
@@
where,
Now, following the formulation of ref. [72] in the novel BGK model, we get
Using eq. (C.112) and the partial derivatives as given in eq. (A.72), eq. (C.110) takes the following form,
For a proper treatment of the left-hand side of the above equation, we are considering an infinitesimal perturbation to the equilibrium distribution function, i.e. with and after linearizing it, we get
where is related to the infinitesimal change of the quark distribution function () by the relation given in eq. (32). After putting the value of in eq. (32), we get as given in eq. (34).
@@ Appendix D Derivation of equation (52)
Appendix E Derivation of equation (58)
In the BGK model, the relativistic Boltzmann transport equation can be written as
With the help of eq. (C), eq. (C.112) and the partial derivatives as given in eq. (D.116), we can express the above equation in the novel BGK model as